OpenStudy (anonymous):

Given the sequence 8, 16, 32, 64, ..., which expression would give the thirteenth term? 8^13 8 · 2^13 8 · 2^12

OpenStudy (anonymous):

a=8 ,r=2 T13=ar^(13-1)=8*2^12

OpenStudy (anonymous):

Where a=first term ,r=common ratio Tn rule : Tn=ar^(n-1) ___ Note :This is a geo.sequence .

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

{Yw} .